Can you PLEASE help me with this homework problem?? I appreciate any help you can give!
Table I
White Black Hispanic Asian Foreign Total
Dept. 1. 4 0 1 1 2 8
Dept. 2. 2 4 1 1 0 8
Dept. 3. 1 2 3 1 1 8
Dept. 4. 4 0 1 1 2 8
Dept. 5. 3 2 1 0 2 8
Total. 14 8 7 4 7 40
Table II
Female Male Total
Black 6 2 8
White 6 8 14
Hispanic 3 4 7
Asian 2 2 4
Foreign 5 2 7
Total 22 18 40

From Table 1: What is the probabilty of choosing a White employee from Dept. 2 and then a Hispanic employee from Dept. 3?
What is the probability of choosing a Foreigner or an Hispanic both from Department 5?
From Table 2: What is the probability of choosing a Black male or a Foreign male?

Answer by stanbon(75887)

Table I
White Black Hispanic Asian Foreign Total
Dept. 1. 4 0 1 1 2 8
Dept. 2. 2 4 1 1 0 8
Dept. 3. 1 2 3 1 1 8
Dept. 4. 4 0 1 1 2 8
Dept. 5. 3 2 1 0 2 8
Total. 14 8 7 4 7 40
Table II
Female Male Total
Black 6 2 8
White 6 8 14
Hispanic 3 4 7
Asian 2 2 4
Foreign 5 2 7
Total 22 18 40
From Table 1: What is the probabilty of choosing a White employee from Dept. 2 and then a Hispanic employee from Dept. 3?
(2/40) * (3/38) = (1/20)*(3/38) = 3/760
-----------------------------------------------------
What is the probability of choosing a Foreigner or an Hispanic both from Department 5?
(1+2)/40 = 3/40
-----------------------
From Table 2: What is the probability of choosing a Black male or a Foreign male?
(2+2)/40 = 1/10
----------------------------
Cheers,
Stan H.